### 25 Mar 2026 v1

@@ Abstract
Single-cell RNA sequencing (scRNA-seq) enables gene expression profiling at cellular resolution, but it is inherently affected by sparsity caused by dropout events, in which expressed genes are recorded as zeros due to technical limitations, such as low mRNA capture efficiency and amplification noise. These artifacts distort gene expression distributions and can compromise downstream analyses. Numerous computational methods have been proposed to address this issue by imputing dropout events and recovering latent transcriptional signals. These methods encompass a wide range of approaches, from traditional statistical models to recently developed deep learning (DL)-based methods. However, their comparative performance remains unclear, as existing benchmarking studies typically evaluate only a limited subset of methods, datasets, and downstream analytical tasks. Here, we present a comprehensive benchmark of 15 scRNA-seq imputation methods spanning 7 methodological categories, including traditional and modern DL-based methods. These methods are evaluated across 30 datasets (26 real and 4 simulated) sourced from 10 experimental protocols and assessed in terms of 6 downstream analytical tasks. Our results show that across the evaluated datasets and analytical tasks, traditional imputation methods, such as model-based, smoothing-based, and low-rank matrix-based methods, generally outperform DL-based methods, such as diffusion-based, GAN-based, GNN-based, and autoencoder-based methods. In addition, strong performance in numerical gene expression recovery does not necessarily translate into improved biological interpretability in downstream analyses, including cell clustering, differential expression analysis, marker gene analysis, trajectory analysis, and cell type annotation. Furthermore, the performance of imputation methods varies substantially across datasets, protocols, and downstream analytical tasks, and no single method consistently outperforms others across all evaluation scenarios. Together, our results provide practical guidance for selecting imputation methods tailored to specific analytical objectives and highlight the importance of task-specific evaluation when assessing imputation performance in scRNA-seq data analysis.
@@
single-cell RNA sequencing, gene expression data, imputation, benchmark
1 Introduction
@@
scRNA-seq imputation methods fall into traditional and deep learning (DL)-based categories [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. Traditional imputation methods typically rely on statistical modeling or similarity-based heuristics [40, 41, 42, 43, 44, 45, 46, 47]. Traditional methods can be broadly categorized into 3 methodological classes, namely model-based, smoothing-based, and low-rank matrix-based methods [16, 19, 36], and are briefly described in Section˜2.3.1. In contrast, DL-based methods rely on representation learning using deep neural networks (DNNs) [48, 49, 50, 51, 52, 53, 54], which is a fundamentally different approach compared to traditional methods. DL-based methods can be broadly categorized into 4 methodological classes, namely diffusion-based, generative adversarial network (GAN)-based, graph neural network (GNN)-based, and autoencoder (AE)-based methods, and are briefly discussed in Section˜2.3.2.
@@
2 Materials and Methods
2.1 Summary of the Data Imputation Benchmarking Framework
Figure 1: Data Imputation Benchmarking Framework.
Fig.˜1 presents an overview of the proposed benchmarking framework, illustrating the complete pipeline from dataset collection and preprocessing to data imputation, downstream tasks, and performance evaluation. Detailed descriptions of dataset collection and preprocessing are provided in Section˜2.2. Data imputation methods, downstream tasks, and evaluation measures are described in Sections˜2.3, 2.4 and 2.5, respectively. Together, this framework enables a systematic and reproducible assessment of scRNA-seq data imputation methods, highlighting their strengths, limitations, and suitability for different analytical scenarios.
2.2 Benchmark Datasets
Table 2: Details of the used scRNA-seq datasets
jurkat Human Jurkat T-cell leukemia cell line scRNA-seq dataset 10x Genomics [^1] https://www.10xgenomics.com/jp/datasets/jurkat-cells-1-standard-1-1-0 $1740\times 13494$ $81.19$ 10x Chromium
pbmc4k Peripheral blood mononuclear cells (PBMCs) from a healthy donor. 10x Genomics [^3] https://www.10xgenomics.com/jp/datasets/4-k-pbm-cs-from-a-healthy-donor-2-standard-2-1-0 $4220\times 16412$ $92.26$ 10x Chromium
simulated_1 - - $2000\times 600$ $30.79$ -
simulated_2 - - $2000\times 600$ $50.62$ -
simulated_3 - - $2000\times 600$ $70.12$ -
simulated_4 - - $2000\times 600$ $89.61$ -
The cells are randomly partitioned into training, validation, and test sets. The training set is used for training the data imputation methods, the validation set is used for early stopping of training iteration of DL-based methods, and the test set is used for evaluating the impuataion methods including running downstream tasks. This dataset splitting strategy prevents data leakage between training and evaluation phases, ensuring a fair assessment of imputation performance, whereas previous benchmarking studies [19, 18, 17] use the same ground truth data for both training and testing, which can lead to overfitting and biased results. The split ratio is $70\,\%$ for training, $10\,\%$ for validation, and $20\,\%$ for testing.
2.3 Data Imputation Methods
2.3.1 Traditional Methods
  • Smoothing-based Methods: Smoothing-based methods leverage the similarity of expression profiles among neighboring cells to reduce noise and preserve biological signals in scRNA-seq data [44, 43, 42], a process typically referred to as smoothing [44, 43, 42]. 3 representative smoothing-based methods are selected in this study, i.e., MAGIC [44], scTsI [43], and AcImpute [42]. MAGIC is designed based on the concept that gene expression profiles can be shared among similar cells to recover missing values, and utilizes Markov processes to impute dropout events [44]. MAGIC first calculates the cell-cell Euclidean distance matrix and constructs a cell-cell affinity matrix using a Gaussian kernel [44]. The affinity matrix is then normalized using row normalization to obtain a Markov transition matrix, which represents the transition probabilities between cells [44]. The process of calculating the Markov matrix is repeated multiple times to reduce noise and keep the biological signal [44]. Finally, the imputed gene expression matrix is obtained by multiplying the final Markov transition matrix with the observed gene expression matrix [44]. scTsI is a 2-stage smoothing-based imputation method that first imputes the zero expression values using the information of neighboring cells and genes, and then adjusts the imputed values using ridge regression [87, 43]. In the first stage, scTsI imputes the zero expression values by calculating an average of the expression values from nearest neighbor cells and nearest neighbor genes [43]. In the second stage, scTsI refines the imputed values by fitting a ridge regression model that predicts the expression value of each gene in each cell based on the expression values of other genes in the same cell and the same gene in other cells [43]. While MAGIC diffuses gene expression values equally across all genes, AcImpute applies different diffusion strengths for highly and lowly expressed genes based on the observation that dropout events are more prevalent in lowly expressed genes [42]. AcImpute first normalizes, selects highly variable genes, and reduces dimensionality using principal component analysis (PCA) [88, 42]. Similar to MAGIC, AcImpute then constructs a cell-cell affinity matrix using $k$-nearest neighbor ($k$-NN)-based adaptive kernel, and obtains a Markov transition matrix through row normalization [42]. In addition, AcImpute calculates the power matrix, a locally averaged diffusion operator, by averaging the normalized matrix over its neighboring cells to capture average gene expression patterns across neighboring cells [42]. Finally, AcImpute combines the Markov transition matrix, the power matrix, and the observed gene expression matrix to create a modified Markov transition matrix [42].
  • Low-rank Matrix-based Methods: Low-rank matrix-based methods are built on the idea that gene expression data often have an underlying simple structure [45, 46, 47]. Specifically, these methods assume that the true gene expression matrix can be well-approximated by a matrix with low rank, meaning that the expression patterns of thousands of genes across many cells can actually be explained by a small number of shared biological factors, such as cell types or cell states [45, 46, 47]. In practice, scRNA-seq data contain a large number of zero values, some of which are caused by dropout events rather than true absence of gene expression [16]. To address this, the observed gene expression matrix can be formulated as $X_{\mathrm{obs}}=X_{\mathrm{true}}+E\in\mathbb{R}^{n\times m}$, where $n$ is the number of cells, $m$ is the number of genes, $X_{\mathrm{true}}$ is the true gene expression matrix to be estimated, and $E$ is a sparse noise matrix that captures dropout events [45, 46, 47]. The objective of low-rank matrix-based methods is to estimate $X_{\mathrm{true}}$ by using the low-rank structure of the gene expression data while accounting for dropout events represented by $E$ [45, 46, 47]. 3 representative low-rank matrix-based methods are selected in this study, i.e., PBLR [45], scLRTC [46], and WEDGE [47]. PBLR explicitly incorporates cell heterogeneity into the imputation process [45]. It first identifies cell subpopulations by constructing multiple cell-cell affinity matrices and applying non-negative matrix factorization (NMF) followed by hierarchical clustering [45]. This step partitions the global expression matrix into more homogeneous submatrices [45]. For each subpopulation-specific submatrix, PBLR performs bounded low-rank matrix recovery, where dropout values are constrained by gene-specific upper bounds estimated from observed expression levels [45]. This bounded formulation prevents unrealistically large imputations and improves recovery accuracy, especially in heterogeneous datasets [45]. scLRTC generalizes matrix-based approaches by modeling scRNA-seq data as a third-order tensor, constructed using cell-cell similarity information [46]. This tensor representation enables simultaneous modeling of gene-gene and cell-cell correlations. scLRTC applies low-rank tensor completion to recover missing values, effectively denoising the data while preserving higher-order structural relationships [46]. By leveraging tensor decomposition rather than simple matrix factorization, scLRTC can better capture complex dependencies in scRNA-seq data and improve downstream analyses such as clustering and trajectory inference [46]. WEDGE addresses dropout by introducing a biased low-rank matrix decomposition framework [47]. Unlike standard matrix factorization methods that ignore zero entries, WEDGE assigns different weights to zero and non-zero elements in the objective function [47]. Non-zero entries are fitted closely to preserve observed expression, while zero entries are softly penalized using a tunable bias parameter, reducing the risk of over-imputation [47]. The model is optimized via alternating non-negative least squares, ensuring biologically meaningful imputed values [47]. This weighted strategy allows WEDGE to robustly recover expression patterns in highly sparse scRNA-seq datasets [47].
2.3.2 DL-based Methods
  • GNN-based Methods: GNN-based methods leverage GNNs to model the relationships between cells and impute dropout events [53]. By propagating information through the graph structure, GNNs can aggregate neighborhood-level features and effectively capture both local and global cellular relationships [94]. In this study, scGNN [53] is included as a representative GNN-based method. scGNN is a hypothesis-free GNN-based method and it integrates 3 iterative multi-modal autoencoders, namely feature AE, graph AE, cluster AE, to model heterogeneous gene expression patterns and aggregate cell-cell relationships [53]. The feature AE receives the regularized gene expression matrix calculated through the left-truncated mixture Gaussian (LTMG) model [95] as input and learns low-dimensional cell representations by minimizing the reconstruction loss between the input and output of the AE [53]. Based on the output of the feature AE, scGNN constructs a cell-cell graph using $k$-NN and feeds it into the graph AE to aggregate neighborhood-level features and learn enhanced cell representations [53]. The cluster AE receives the reconstructed gene expression matrix from the feature AE and an individual encoder is used for each cell cluster to better capture cluster-specific gene expression patterns, which are identified through clustering on the output of the graph AE [53]. The reconstructed gene expression matrices from an individual encoder of the cluster AE are concatenated, and fed into the feature AE and graph AE in the next iteration [53]. This iterative process continues until convergence, and the final reconstructed gene expression matrix from the feature AE is used as the imputed gene expression matrix [53].
2.4 Downstream Tasks
The quality of imputed scRNA-seq data directly influences the reliability and performance of computational models in downstream tasks [17, 19, 18]. Inaccurate or biased imputation can distort underlying biological signals, leading to misleading conclusions [17, 19, 18]. Therefore, a comprehensive evaluation of imputation methods must assess not only numerical recovery of gene expression values but also their impact on biologically meaningful downstream tasks [17, 19, 18]. In this section, as illustrated in Fig.˜2, we describe 6 distinct downstream tasks used to benchmark imputation methods, namely numerical gene expression recovery, cell clustering, DE analysis, marker gene analysis, trajectory analysis, and cell type annotation.
  • Numerical Gene Expression Recovery: Numerical gene expression recovery can be formulated as a regression task, in which the objective is to predict true gene expression values from sparsely observed scRNA-seq data affected by dropout events [17]. In this setting, the input consists of corrupted expression matrices where zero or near-zero values arise due to technical dropouts, while the target outputs correspond to the original, uncorrupted gene expression values [36]. Ground truth data are obtained from real datasets where artificial dropout is introduced in a controlled manner [36, 17, 19]. Imputation models are trained by learning a mapping from the observed sparse data to the complete expression space, and performance is quantitatively evaluated using regression-based error metrics such as mean squared error (MSE) and mean absolute error (MAE) computed at the gene, cell, or matrix level [17, 18].
  • Cell Clustering: Cell clustering can be formulated as an unsupervised learning problem, where each cell is treated as an individual sample represented by a high-dimensional gene expression vector [55]. The objective is to group similar cells into clusters based on their expression profiles [55]. Since scRNA-seq data is inherently high-dimensional, with thousands of genes measured per cell, dimensionality reduction techniques, such as PCA [88], t-distributed stochastic neighbor embedding (t-SNE) [99], and uniform manifold approximation and projection (UMAP) [100], are often applied prior to clustering to reduce noise and improve computational efficiency [2]. Cell clustering is typically performed as an initial downstream task and serves as a foundation for subsequent downstream tasks, such as marker gene analysis and cell type annotation [55, 2]. Meaningful clusters enable the identification of distinct cell populations and cellular states, whereas inaccurate clustering can lead to misleading biological interpretations [2]. In this benchmark, dimensionality reduction is performed using PCA [88], and subsequently apply the Leiden algorithm [21] for cell clustering. The Leiden algorithm is a graph-based community detection algorithm that operates on a cell-cell similarity graph constructed from the scRNA-seq data and partitions cells into clusters by optimizing modularity, a quality function that measures the density of edges within clusters compared to edges between clusters [21].
  • DE analysis: DE analysis can be formulated as a feature selection problem, in which the objective is to identify genes that exhibit statistically significant expression differences between conditions, such as disease versus healthy groups or treatment versus control groups [2, 30]. This is typically achieved by testing the null hypothesis that the expression levels of a given gene are identical between the 2 groups [101]. DE analysis enables the identification of genes associated with specific biological processes or disease states and represents a core downstream task for linking gene expression changes to underlying biological phenomena [2]. Accurate identification of differentially expressed genes (DEGs) is therefore crucial for understanding molecular mechanisms related to disease progression or treatment response [2]. In this benchmark, 2 common DE analysis methods are used, namely MAST [102], which is a statistical framework using a hurdle model to account for the bimodal distribution of scRNA-seq data [102, 19], and the Wilcoxon rank-sum test [103, 104, 105], in order to evaluate the performance and robustness of imputation methods across multiple DE analysis approaches [19].
  • Marker Gene Analysis: Marker gene analysis is an application of DE analysis, and can be formulated as a feature selection problem, in which the objective is to identify genes that best represent each cluster of cells [2, 28]. This task is typically performed in 2 steps. First, DE analysis is conducted to identify genes whose expression levels significantly differ between a given cluster and the remaining clusters [2, 28]. Second, genes are ranked based on log-fold change (LFC) or test statistics derived from DE analysis, and the top-ranked genes are selected as marker genes for each cluster [28]. Marker gene analysis plays a crucial role in the interpretation of cell clusters, as marker genes provide insights into the biological functions and identities of distinct cell populations [28, 2]. Accurate identification of marker genes enables reliable interpretation of the biological significance of cell clusters, and a deeper understanding of underlying cellular heterogeneity [28, 2]. In this benchmark, marker gene analysis is performed using the Wilcoxon rank-sum test [103, 104, 105]-based DE analysis between each cluster and the remaining clusters.
  • Cell Type Annotation: Cell type annotation can be formulated as a supervised multi-class classification problem, in which inputs are gene expression vectors of each cell, and outputs are corresponding cell type labels [2, 56, 55]. The objective of this task is to learn a mapping from gene expression vectors to cell type labels based on known cell type labels [56]. Cell types are predefined at different levels of granularity, such as broad cell types, e.g., T cells, B cells, and monocytes, or fine-grained cell subtypes, e.g., CD4+ T cells, CD8+ T cells, and regulatory T cells [56, 108]. A typical approach to perform cell type annotation is to compare the scRNA-seq data with previously annotated reference datasets using classification models [56]. In this approach, a classifier is trained on a reference dataset to learn the mapping from gene expression vectors to cell type labels, and is subsequently used to predict cell type labels for new scRNA-seq data [56]. This task is essential for the interpretation of scRNA-seq data, as it provides biological context for cells and clusters identified in the data [56, 55, 16]. Robust cell type annotation enables researchers to better understand cellular heterogeneity and the functional roles of different cell types in biological processes [2, 56]. In this benchmark, scGPT [109], a foundation model for scRNA-seq data which supports cell type annotation, and 1D convolutional neural network (1D-CNN) are used to evaluate the performance of cell type annotation across different imputation methods and scRNA-seq datasets. For scGPT, the pretrained model released by the authors [^1] https://github.com/bowang-lab/scGPT, which is trained on 33 million human cells, is used, and for 1D-CNN, the model is trained on the training set of each dataset.
2.5 Evaluation Measures
  • Numerical Gene Expression Recovery: The evaluation of numerical gene expression recovery can be performed by directly comparing imputed gene expression values with ground truth values or by comparing with corresponding bulk RNA-seq data [17, 19, 18]. 3 distinct evaluation measures are used to directly evaluate the numerical gene expression recovery performance of all 15 imputation methods, namely MAE, median absolute error (MedAE), log normalized difference (LND), and MSE. Each measure is computed by comparing the imputed gene expression values with the ground truth expression values. MAE is computed as the average of the absolute differences between imputed and ground truth values, which provides a more direct measure of average error [17]. MedAE is computed as the median of the absolute differences between imputed and ground truth values, which provides a similar measure as MAE without outliers [17]. LND is calculated as the log-transformed difference between imputed and ground truth values, which allows assessment of over- or under-imputation [17]. MSE is calculated as the average of the squared differences between imputed and ground truth values, which provides a measure of overall error magnitude [17, 18]. In addition to comparison with ground truth values, comparison with matched bulk RNA-seq data is performed to evaluate the performance of imputation methods in recovering gene expression patterns [19] with 2 distinct evaluation measures, namely pseudo-bulk correlation coefficient (PCC) and median correlation coefficient (MCC). PCC measures the correlation between pseudo-bulk expression values, calculated by averaging imputed gene expression values across all cells, and bulk RNA-seq expression values [19]. MCC measures the median correlation between imputed gene expression values of individual cells and bulk RNA-seq expression values [19]. Both measures are calculated using Spearman’s rank correlation coefficient (SCC) [110].
  • Cell Clustering: 4 distinct evaluation measures are used to evaluate cell clustering performance, namely adjusted rand index (ARI) [111], normalized mutual information (NMI), purity, and silhouette coefficient (SC) [17, 18, 19]. ARI measures the similarity between 2 clustering results by considering all pairs of cells and counting pairs that are assigned in the same or different clusters in the clustering results of imputed data and ground truth data [17, 18, 19]. NMI measures the mutual dependence between the clustering results of imputed data and ground truth data [17]. Purity measures clustering quality by quantifying how homogeneous each predicted cluster is with respect to ground truth labels [17]. SC measures clustering quality by quantifying cohesion within clusters and separation between clusters [17, 18].
  • DE analysis: 2 distinct evaluation measures are used to evaluate DE analysis performance, namely intersection over union (IoU) and false positive DEG (FPDEG). IoU measures the overlap between the sets of genes in the 2 different groups, e.g., DEGs identified from imputed scRNA-seq data and from the corresponding bulk RNA-seq data. FPDEG measures the number of genes identified as DEGs that are not true DEGs.
  • Trajectory Analysis: 2 distinct evaluation metrics are used to evaluate trajectory analysis performance, namely pseudo-temporal ordering score (POS) [107, 18] and Kendall’s rank correlation coefficient (KRCC) [112, 18]. POS is calculated by summing scores that characterize how well the inferred cell ordering matches the expected ordering based on external information. KRCC is computed to measure the correlation between the inferred pseudotime values and the true cell development labels.
    $f(x)=\begin{cases}\text{Pseudo-temporal Ordering Score (POS)}=\sum_{i=1}^{n-1}\sum_{j>i}^{n}g(i,j)\\ \text{Kendall's Rank Correlation Coefficient (KRCC)}=\frac{4C}{n(n-1)}-1\end{cases}$ (4)
  • Cell Type Annotation: 4 distinct evaluation measures are used to evaluate cell type annotation performance, namely macro accuracy (ACC), macro precision (PR), macro recall (RC), and macro F1 score (F1). Each measure is computed using a macro-averaging approach to ensure equal weighting for all cells irrespective of their types. ACC is calculated as the average of individual accuracy scores across all cells. For a single cell, accuracy is computed as the ratio of correctly predicted samples to the total samples in that cluster. PR is calculated as the average of the ratio of true positives to total predicted positives for each cell, with single-cell PR computed as true positives divided by predicted positives. RC is the average of the ratio of true positives to total actual positives for each cell, with single-cell RC computed as true positives divided by actual positives. F1 is the harmonic mean of PR and RC, calculated across all cells. For individual cells, the F1 is computed as the harmonic mean of that cell’s PR and RC.
2.6 Experimental Setup
@@ 3.1 Numerical Gene Expression Recovery
Fig.˜3 represents the distribution of LND values between imputed and ground truth expression values for 15 imputation methods in terms of 26 real and 4 simulated datasets. The width of each violin plot represents the density of LND values, and the box plot shows the median and interquartile range (IQR) of LND values. $\mathrm{LND}=0$ indicates that all imputed expression values are identical to the ground truth values, $\mathrm{LND}>0$ indicates over-imputation, where the imputed values are greater than the ground truth values, and $\mathrm{LND}<0$ indicates under-imputation, i.e., the imputed values are less than the ground truth values.
@@
Figs. 4a and b represent the box plots for MAE and MedAE between imputed and ground truth expression values for 15 imputation methods in terms of 26 real and 4 simulated datasets. MAE shows the overall performance of the methods considering outliers, while MedAE shows the overall performance of the methods without outliers. Lower MAE and MedAE values indicate better recovery performance, as they suggest that the imputed expression values are closer to the ground truth values.
@@
Fig.˜4c shows total MAE and total MedAE across all datasets for each imputation method. WEDGE shows the best MAE across 5 protocols, namely 10x Chromium, CEL-seq2, SMART-seq2, SMART-seq, and Drop-seq. Similarly, WEDGE shows the best MedAE across 3 protocols, namely 10x Chromium, CEL-seq2, and Drop-seq. On the other hand, scIDPMs and scIGANs exhibit the highest MAE and MedAE as their values significantly exceed the masked baseline for all protocols.
Figs. 4d and e show PCC between pseudo-bulk and the corresponding bulk RNA-seq data for 15 imputation methods in terms of 2 cell line datasets, namely sc_10x_5cl and encode_fluidigm_5cl. A higher PCC indicates that pseudo-bulk data is highly correlated with the corresponding bulk RNA-seq data. The comparison of PCC across the 15 methods shows that AcImpute, WEDGE, and CPARI achieve the best performance. In addition, 10 methods, namely PbImpute, scImpute, scTsI, MAGIC, PBLR, scLRTC, scIDPMs, stDiff, scGNN, and Bubble, show moderate performance. On the other hand, scMultiGAN and scIGANs show the worst performance.
Fig.˜4f shows MCC between imputed scRNA-seq data and the corresponding bulk RNA-seq data at the cell line level for 15 imputation methods in terms of 2 cell line datasets, namely sc_10x_5cl and encode_fluidigm_5cl. A higher MCC indicates that imputed scRNA-seq data is highly correlated with the corresponding bulk RNA-seq data. The comparison of MCC across the 15 methods shows that MAGIC and WEDGE achieve the best performance. In addition, 10 methods, namely PbImpute, scImpute, AcImpute, scTsI, PBLR, scLRTC, stDiff, scGNN, CPARI, and Bubble, show moderate performance. On the other hand, scIDPMs, scMultiGAN, and scIGANs show the worst performance.

### 13 Apr 2026 v2

@@ Abstract
Background: Single-cell RNA sequencing (scRNA-seq) enables gene expression profiling at cellular resolution but is inherently affected by sparsity caused by dropout events, where expressed genes are recorded as zeros due to technical limitations. These artifacts distort gene expression distributions and compromise downstream analyses. Numerous imputation methods have been proposed to recover latent transcriptional signals. These methods range from traditional statistical models to deep learning (DL)-based methods. However, their comparative performance remains unclear, as existing benchmarks evaluate only a limited subset of methods, datasets, and downstream analyses.
Results: We present a comprehensive benchmark of 15 scRNA-seq imputation methods spanning 7 methodological categories, including traditional and DL-based methods. Methods are evaluated across 30 datasets from 10 experimental protocols on 6 downstream analyses. Results show that traditional methods, such as model-based, smoothing-based, and low-rank matrix-based methods, generally outperform DL-based methods, including diffusion-based, generative adversarial network-based, graph neural network-based, and autoencoder-based methods. In addition, strong performance in numerical gene expression recovery does not necessarily translate into improved biological interpretability in downstream analyses, including cell clustering, differential expression analysis, marker gene analysis, trajectory analysis, and cell type annotation. Furthermore, method performance varies substantially across datasets, protocols, and downstream analyses, with no single method consistently outperforming others.
Conclusions: Our findings provide practical guidance for selecting imputation methods tailored to specific analytical objectives and underscore the importance of task-specific evaluation when assessing imputation performance in scRNA-seq data analysis.
@@
single-cell RNA sequencing, gene expression, imputation, benchmark
1 Background
@@
scRNA-seq imputation methods fall into traditional and deep learning (DL)-based categories [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. Traditional imputation methods typically rely on statistical modeling or similarity-based heuristics [40, 41, 42, 43, 44, 45, 46, 47]. Traditional methods can be broadly categorized into 3 methodological classes, namely model-based, smoothing-based, and low-rank matrix-based methods [16, 19, 36], and are briefly described in Section˜5.3.1. In contrast, DL-based methods rely on representation learning using deep neural networks (DNNs) [48, 49, 50, 51, 52, 53, 54], which is a fundamentally different approach compared to traditional methods. DL-based methods can be broadly categorized into 4 methodological classes, namely diffusion-based, generative adversarial network (GAN)-based, graph neural network (GNN)-based, and autoencoder (AE)-based methods, and are briefly discussed in Section˜5.3.2.
@@ 2.1 Numerical Gene Expression Recovery
Fig.˜1 represents the distribution of log normalized difference (LND) values between imputed and ground truth expression values for 15 imputation methods in terms of 26 real and 4 simulated datasets. The width of each violin plot represents the density of LND values, and the box plot shows the median and interquartile range (IQR) of LND values. $\mathrm{LND}=0$ indicates that all imputed expression values are identical to the ground truth values, $\mathrm{LND}>0$ indicates over-imputation, where the imputed values are greater than the ground truth values, and $\mathrm{LND}<0$ indicates under-imputation, i.e., the imputed values are less than the ground truth values.
@@
Figs. 2a and b represent the box plots for mean absolute error (MAE) and median absolute error (MedAE) between imputed and ground truth expression values for 15 imputation methods in terms of 26 real and 4 simulated datasets. MAE shows the overall performance of the methods considering outliers, while MedAE shows the overall performance of the methods without outliers. Lower MAE and MedAE values indicate better recovery performance, as they suggest that the imputed expression values are closer to the ground truth values.
@@
Fig.˜2c shows total MAE and total MedAE across all datasets for each imputation method. WEDGE shows the best MAE across 5 protocols, namely 10x Chromium, CEL-seq2, SMART-seq2, SMART-seq, and Drop-seq. Similarly, WEDGE shows the best MedAE across 3 protocols, namely 10x Chromium, CEL-seq2, and Drop-seq. On the other hand, scIDPMs and scIGANs exhibit the highest MAE and MedAE as their values significantly exceed the masked baseline for all protocols.
Figs. 2d and e show pseudo-bulk correlation coefficient (PCC) between pseudo-bulk and the corresponding bulk RNA-seq data for 15 imputation methods in terms of 2 cell line datasets, namely sc_10x_5cl and encode_fluidigm_5cl. A higher PCC indicates that pseudo-bulk data is highly correlated with the corresponding bulk RNA-seq data. The comparison of PCC across the 15 methods shows that AcImpute, WEDGE, and CPARI achieve the best performance. In addition, 10 methods, namely PbImpute, scImpute, scTsI, MAGIC, PBLR, scLRTC, scIDPMs, stDiff, scGNN, and Bubble, show moderate performance. On the other hand, scMultiGAN and scIGANs show the worst performance.
Fig.˜2f shows median correlation coefficient (MCC) between imputed scRNA-seq data and the corresponding bulk RNA-seq data at the cell line level for 15 imputation methods in terms of 2 cell line datasets, namely sc_10x_5cl and encode_fluidigm_5cl. A higher MCC indicates that imputed scRNA-seq data is highly correlated with the corresponding bulk RNA-seq data. The comparison of MCC across the 15 methods shows that MAGIC and WEDGE achieve the best performance. In addition, 10 methods, namely PbImpute, scImpute, AcImpute, scTsI, PBLR, scLRTC, stDiff, scGNN, CPARI, and Bubble, show moderate performance. On the other hand, scIDPMs, scMultiGAN, and scIGANs show the worst performance.